Use an algebraic approach to solve the problem.
If a ring costs a jeweler $180, at what price should it be sold to yield a profit of 50% on the selling price?

Answer by stanbon(75887)

If a ring costs a jeweler $180, at what price should it be sold to yield a profit of 50% on the selling price?
------
Price = 180 + 0.5*180 = 1.5*180 = $270
========================================
Cheers,
Stan H.
==============